OpenStudy (anonymous):

Once the support structure is built, it is covered with a triangular piece of plywood. The area of the piece of wood required to cover triangle ABC is ____ cm2. (Enter only the number.) Hint: Look back at the labels on the original drawing. Add the segments to find the base of the triangle, and use the given height in the formula for area of the triangle, A = (1/2)(b)(h).

OpenStudy (anonymous):

Once the support structure is built, it is covered with a triangular piece of plywood. The area of the piece of wood required to cover triangle ABC is ____ cm2. (Enter only the number.) Hint: Look back at the labels on the original drawing. Add the segments to find the base of the triangle, and use the given height in the formula for area of the triangle, A = (1/2)(b)(h).

OpenStudy (phi):

The hint practically gives you the answer. Add up the 4 numbers at the bottom to get the distance from B to C. They give you the height AB= 50 cm. The area is just what they say in the hint.

OpenStudy (anonymous):

Phi, so is this right? A=(1/2) (150)(50) = \[3750 cm2\]